OpenStudy (anonymous):

Find the derivative of f(x) = 6/x at x = -2.

OpenStudy (anonymous):

@agent0smith

OpenStudy (agent0smith):

Do it the same way as the last, f'(x) = (f(x+h) - f(x))/h this time, f(x+h) = 6/(x+h) f'(x) = ( 6/(x+h) -6/x)/h You'll have to simplify those fractions, though...

OpenStudy (anonymous):

oh poop lolol

OpenStudy (agent0smith):

Yeah I don't really want to try it w/o the equation editor....

OpenStudy (agent0smith):

easier to read: http://www.wolframalpha.com/input/?i=simplify+6%2F%28x%2Bh%29+-6%2Fx

OpenStudy (agent0smith):

Now if you divide that result by h, then plug in h=0, you'll have the derivative

OpenStudy (e.mccormick):

http://www.codecogs.com/latex/eqneditor.php You can make links or images from that. Just be sure to make the background white.

OpenStudy (anonymous):

-6 ?

OpenStudy (e.mccormick):

I think you made a small mistake in simplifying.

OpenStudy (agent0smith):

http://www.wolframalpha.com/input/?i=simplify+%28+6%2F%28x%2Bh%29+-6%2Fx%29%2Fh now plug in h=0 into either of those simplified results. (the derivative is really the limit as h approaches zero)

OpenStudy (anonymous):

wait 6? lol

OpenStudy (agent0smith):

-6/(hx + x^2) plug in h=0

OpenStudy (anonymous):

i gots -6/x^2

OpenStudy (e.mccormick):

AH HA! that looks better.

OpenStudy (agent0smith):

Yes :)

OpenStudy (e.mccormick):

Now evaluate at f'(-2)

OpenStudy (anonymous):

so would i put x=-2^2

OpenStudy (agent0smith):

f'(x) = -6/x^2 plug in x=-2

OpenStudy (anonymous):

-6/-2^2

OpenStudy (e.mccormick):

well, x=-2, not -2^2. It will be squared, but not before you plug it in.

OpenStudy (e.mccormick):

Yah, -6/(-2)^2

OpenStudy (anonymous):

i got -1.5

OpenStudy (agent0smith):

Correct :)

OpenStudy (e.mccormick):

Now you may not have known what my whole tangent line thing was. This is what it would look like in this case: https://www.desmos.com/calculator/jvebfrvtvj I doubt you are doing those yet, but just thought I would show how the whole knowing point-slope formula comes into play later as you get into calculus.

OpenStudy (anonymous):

so 3/2?

OpenStudy (anonymous):

well-3/2

OpenStudy (e.mccormick):

Yah. -3/2, whch is -1.5.

OpenStudy (anonymous):

@e.mccormick and @agent0smith THANK YOU SO MUCH GUYS!!♥

OpenStudy (agent0smith):

Welcome ♥

OpenStudy (e.mccormick):

np. Have fun. And I hope you can see how calculus depends on algebra, so all that learning you did before this will come into use later. Well, calculus is used in medicine, physics, engineering, and so on, so everything it takes to learn calculus comes into play in doing so many other things. Well worth understanding it for that.